OpenStudy (anonymous):

need to find the equation of the line having the given slope and containing the given point m=5, (8,4)

OpenStudy (anonymous):

Are you familiar with the point slope formula? It basically does exactly what you're asking for here..

OpenStudy (anonymous):

yes, but for some reason I blank looking at it

OpenStudy (anonymous):

Ok, so lets say you are given a slope (m) and a point (h,k) Then the equation for the line going through that point having that slope is: \[y-k = m(x-h)\] That's it. In this case (h,k) is (8,4) and m is 5. Find & replace.

OpenStudy (anonymous):

Give it a try?

OpenStudy (anonymous):

im working at it but im going to warn u I really suck at this...lol i draw blanks

OpenStudy (anonymous):

There's nothing to do. What is h? what is k? what is m?

OpenStudy (anonymous):

i know im just writing out

OpenStudy (anonymous):

Well can you tell me really fast what h is?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

So instead of \[y-k = m(x-h)\] we have \[y-k = m(x-8)\] What is k?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So we have: \[y-4 = m(x-8)\] What is m?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

So we have:\[y-4 =5(x-8)\] And we're finished.

OpenStudy (anonymous):

see i know its easy but for some reason my brain doesnt connect with math. Thank u so much

OpenStudy (anonymous):

When they say 'h' is '8' it literally means you can take one, and put in the other instead and it means the same thing.

OpenStudy (anonymous):

i need to keep u around

OpenStudy (anonymous):

u make sense

OpenStudy (anonymous):

Most people just get themselves all stressed out when it comes to math. There's nothing tricky about it. They try to dress it up with funny symbols, but once you know what it means it's really very straight forward.

OpenStudy (anonymous):

Don't panic. You can do this. You are good at it, you just confuse yourself or put up roadblocks.

OpenStudy (anonymous):

very true. I have always stressed out when it came to math because its the only thing I copuld never do for whatever reason

OpenStudy (anonymous):

All math is just ordered thinking. Has almost nothing to do with numbers.

OpenStudy (anonymous):

I could do it at work but I cant do it for school

OpenStudy (anonymous):

You can. You have to believe that you can do it, and that it's not as hard as you think. =)

OpenStudy (anonymous):

95% of peoples issues with math have to do with social and environmental training and stress. Almost nothing to do with aptitude.

OpenStudy (anonymous):

I believe I can do it but I know I cant...lol I get stuck on stupid

OpenStudy (anonymous):

ever since I can remember I could never do math

OpenStudy (anonymous):

im jinxed when it comes to Math....lol :)

OpenStudy (anonymous):

That's the problem right there though. Not that you cannot do it, but that you believe you cannot. You must change the way you think about math. It is not a chore. It's not a punishment. It's just a way of thinking about things (everything) and you honestly do it all the time without realizing it.

OpenStudy (anonymous):

I do it at work everyday with no problem but for some reason because I HAVE to do it for school, i freak out. You are right about that

OpenStudy (anonymous):

So lets look at another problem and try not to freak out. Or if you don't have any, I can make some up for you. You're studying lines?

OpenStudy (anonymous):

yes, im studying lines. Let me see, I may have another

OpenStudy (anonymous):

Even if it's similar it's good to practice what you know.

OpenStudy (anonymous):

for that last one though I need the y= part

OpenStudy (anonymous):

really? Ok then. What do you need to do first?

OpenStudy (anonymous):

Actually hang on.

OpenStudy (anonymous):

We can make a new formula! That will make this easier the next time you have to do it.

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

So lets see.. we started with a slope and a point. Lets call the slope m and the point (h,k) again. What was the point slope formula?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

That's the slope intercept formula.

OpenStudy (anonymous):

Same line, but different arrangement that makes some things easier to see.

OpenStudy (anonymous):

thats what it asks me to put the answer in though

OpenStudy (anonymous):

That's fine.. That's where we want to end up. I'm saying where did we start?

OpenStudy (anonymous):

What was the formula I was having you plug in h,k and m into?

OpenStudy (anonymous):

y-k=m(x-h)

OpenStudy (anonymous):

Indeed. So now we need to have the y by itself to find y = mx + b form. What do we need to do?

OpenStudy (anonymous):

i know we need to subtract something from both sides

OpenStudy (anonymous):

That's right. (or add perhaps). What can we add or subtract from both sides to get the y by itself?

OpenStudy (anonymous):

What specifically do we need to get rid of on the left hand side?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Right. So what do we do to get rid of the -k ?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

add what?

OpenStudy (anonymous):

k (4)

OpenStudy (anonymous):

We're just dealing with h , k and m now because then we won't have to re-do all this when we have another problem and they tell us h,k is 20,30 etc.

OpenStudy (anonymous):

So if we add a k to both sides, what will we get?

OpenStudy (anonymous):

What will we get on the left side (the whole reason we did it)?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Right, and on the right side? We still have what we had before +k right?

OpenStudy (anonymous):

So rewrite the equation?

OpenStudy (anonymous):

y=5(x-8)

OpenStudy (anonymous):

no no. Where did those numbers come from?

OpenStudy (anonymous):

Just h, k, and m.

OpenStudy (anonymous):

y = m(x-h) + k

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

o yes sorry

OpenStudy (anonymous):

So now we just have to simplify the m(x-h) bit into the form we want.

OpenStudy (anonymous):

Can you distribute that m to each of the terms?

OpenStudy (anonymous):

so am I plugging in numbers now?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Wait till we get the form we want. Then plug in the numbers after all the algebra is done.

OpenStudy (anonymous):

Distribute, m(x - h) = ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

m(x-h) = mx - mh

OpenStudy (anonymous):

So we have: y = mx -(mh - k) = mx + b So b = -(mh - k)

OpenStudy (anonymous):

i couldnt type it fast enough

OpenStudy (anonymous):

So now we can plug in any numbers you have for m, h, and k to find your equation.

OpenStudy (anonymous):

y=-4(5x-8)? or something close

OpenStudy (anonymous):

\[y = 5x -(5(8) - 4)\] \[\implies y = 5x -(40 - 4)\]\[\implies y = 5x - 36\]

OpenStudy (anonymous):

wouldnt it be y=5x+36